OpenStudy (milrawr):

The expected pop of a certain bird grows according to the model p(t)=(500)/(1+83.33e^-0.162t) When will the population reach 300?

OpenStudy (irishboy123):

you want to solve: \[p(t) = \dfrac{500}{1+83.33e^{-0.162t}} \color{green}{=300} \] that means you have to say that: \[\dfrac{500}{300} =1+83.33e^{-0.162t} \] Such that: \[83.33~ e^{-0.162t} = \dfrac{500}{300} - 1 \] and so on..... :-)) So process the algebra and take some logs.

OpenStudy (milrawr):

@IrishBoy123 how come the 500 just went over to the other side how'd you know that

OpenStudy (irishboy123):

You can do it many ways: \[\dfrac{500}{1+83.33e^{-0.162t}} \color{green}{=300}\] \[\dfrac{1+83.33e^{-0.162t}}{500} =\dfrac{1}{300}\] \[1+83.33e^{-0.162t} =\dfrac{500}{300}\] \[83.33e^{-0.162t} =\dfrac{500}{300} - 1\] \[e^{-0.162t} =\dfrac{1}{83.33} \left( \dfrac{500}{300} - 1 \right)\] \[-0.162t =\ln \left( \dfrac{1}{83.33} \left( \dfrac{500}{300} - 1 \right) \right)\]

OpenStudy (milrawr):

@IrishBoy123 I'm really confused on how to get those possibilities ^ I can't seem to even get 1

OpenStudy (irishboy123):

post your own approach ! we can work through that people approach these things in different ways. Most of the times, i reckon :-)

OpenStudy (milrawr):

how do i write the exponents?

OpenStudy (irishboy123):

that last thing i posted is merely a question of your posting some numbers into a calculator This \[-0.162t =\ln \left( \dfrac{1}{83.33} \left( \dfrac{500}{300} - 1 \right) \right)\] Followed by: \[t =-\dfrac{1}{0.162} \ln \left( \dfrac{1}{83.33} \left( \dfrac{500}{300} - 1 \right) \right)\] there could easily be a typo in there as I have just latexed this on OS :))